### v4

# SEMF: Supervised Expectation-Maximization Framework for Predicting Intervals

Abstract
This work introduces the Supervised Expectation-Maximization Framework (SEMF), a versatile and model-agnostic approach for generating prediction intervals with any ML model. SEMF extends the Expectation-Maximization algorithm, traditionally used in unsupervised learning, to a supervised context, leveraging latent variable modeling for uncertainty estimation. Through extensive empirical evaluation of diverse simulated distributions and 11 real-world tabular datasets, SEMF consistently produces narrower prediction intervals while maintaining the desired coverage probability, outperforming traditional quantile regression methods. Furthermore, without using the quantile (pinball) loss, SEMF allows point predictors, including gradient-boosted trees and neural networks, to be calibrated with conformal quantile regression. The results indicate that SEMF enhances uncertainty quantification under diverse data distributions and is particularly effective for models that otherwise struggle with inherent uncertainty representation. [^1] Code: https://github.com/Unco3892/semf-paper
1 Introduction
In the evolving field of machine learning (ML), the quest for models able to predict outcomes while quantifying the uncertainty of their predictions is critical. The ability to estimate prediction uncertainty is particularly vital in high-stakes domains such as healthcare [1], finance [2], and autonomous systems [3], where prediction-based decisions have important consequences. Traditional approaches have primarily focused on point estimates, with little to no insight into prediction reliability. This limitation underscores the need for frameworks that can generate both precise point predictions and robust prediction intervals. Such intervals provide a range within which the true outcome is expected to lie with a fixed probability, offering a finer understanding of prediction uncertainty. This need has spurred research into methodologies that extend beyond point estimation to include uncertainty quantification, thereby enabling more informed decision-making in applications reliant on predictive modeling [4].
In this paper, we introduce the Supervised Expectation-Maximization Framework (SEMF) based on the Expectation-Maximization (EM) algorithm [5]. Traditionally recognized as a clustering technique, EM is used for supervised learning in SEMF, allowing for both point estimates and prediction intervals using any ML model (model-agnostic), with a particular focus on uncertainty quantification. This paper details the methodology behind the framework and proposes a training algorithm based on Monte Carlo (MC) sampling, also used in variational inference for Variational Auto-Encoders (VAEs) [6, 7]. The EM algorithm has already been used for supervised learning tasks using specific models [8, 9, 10], minimal supervision [11], and using labeled and unlabeled data in semi-supervised settings for both single and multiple modalities [12, 13]. Our method, SEMF, differs from prominent supervised EM approaches such as ghahramani1993em [8], which focus on point prediction using Gaussian Mixture Models (GMMs). Furthermore, SEMF operates in a frequentist paradigm, directly maximizing the likelihood function through iterative EM steps without integrating over posterior distributions. Although SEMF can be extended to a Bayesian framework as its likelihood component, this extension lies beyond the scope of this paper. Finally, SEMF generates representations for latent modalities through specialized models, holding potential for multi-modal data applications, though multi-modal settings are left for future exploration.
The remainder of this paper is organized as follows: Section 2 details the theory and methodology of SEMF. Section 3 reviews related work in latent representation learning and uncertainty estimation. Section 4 describes the experimental setup, including synthetic and benchmark datasets, as well as evaluation metrics. Section 5 discusses the results, demonstrating the efficacy of SEMF. Lastly, Section 6 concludes the paper, and Section 7 outlines the limitations and potential research directions.
2 Method
This section presents the founding principles of SEMF from its parameters, training, and inference procedure with, at its core, the EM algorithm. Invented to maximize the model likelihood, it builds a sequence of parameters that guarantee an increase in the log-likelihood [14] by iterating between the Expectation (E) and the Maximization (M) steps. In the E-step, one computes
[…]
where $p^{\prime}$ stands for the current estimates, $\log p(x,z)$ is the log-likelihood of the complete observation $(x,z)$, and $z$ is a latent variable. The M-step updates the current estimates with the arguments maximizing the $Q$-function: $p^{\prime}\leftarrow\arg\max_{p}Q(p|p^{\prime})$ The sequence is repeated until convergence.
2.1 Problem Scenario
Let $x=(x_{1},x_{2},\ldots,x_{K})$ denote $K$ inputs and the output be $y$. For simplicity, we limit $y$ to be continuous numerical, although this assumption could be relaxed to discrete or categorical without loss of generality. Component $x_{k}$ is a source: a modality, a single or group of variables, or an unstructured input such as an image or text. For clarity, we limit to $K=2$, where $x_{1}$ and $x_{2}$ are single variables.
[…]
The contribution to the log-likelihood of a complete observation $(y,z,x)$ is $\log p(y,z|x)=\log p(y|z)+\log p(z|x)$. In the E-step, we compute
[…]
Eq. 2 can be estimated by MC sampling. Since sampling from $p^{\prime}(z|y,x)$ can be inefficient, we rather rely on the decomposition $p^{\prime}(z|y,x)=p^{\prime}(y|z)p^{\prime}(z|x)/p^{\prime}(y|x)$. Thus, we sample $z_{r}$ from $p^{\prime}(z|x)$, $r=1,\ldots,R$, and, setting $w_{r}=p^{\prime}(y|z_{r})/\sum_{t}p^{\prime}(y|z_{t})$, approximate the right-hand side term of Eq. 2
[…]
2.2 Objective Function
Adapting Eq. 3 for the observed data $\{(y_{i},x_{i})\}_{i=1}^{N}$, the overall loss function, $\mathcal{L}$, is
$\mathcal{L}(\phi,\theta)=-\sum_{i=1}^{N}\sum_{r=1}^{R}\left\{\log p_{\phi}(z_{i,r}|x_{i})+\log p_{\theta}(y_{i}|z_{i,r})\right\}w_{i,r},$ (4)
where the models of $p(y|z)$ and $p(z|x)$ inherit parameters $\theta$ and $\phi$, respectively. The weights are
$w_{i,r}=\frac{p_{\theta^{\prime}}(y_{i}|z_{i,r})}{\sum_{t=1}^{R}p_{\theta^{\prime}}(y_{i}|z_{i,t})}$ (5)
where $z_{i},r\stackrel{{\scriptstyle ind.}}{{\sim}}p_{\phi^{\prime}}(z|x_{i})$, $r=1,\ldots,R$.
Eq. 4 shows that ${\cal L}$ is a sum of losses associated with the encoder model, $p_{\phi}$, for each source, and the decoder model, $p_{\theta}$, from the latent variables to the output. $p_{\theta}$ and $p_{\phi}$ are referred to as encoder and decoder in the spirit of auto-encoder models. At each M-step, ${\cal L}$ is minimized with respect to $\theta$ and $\phi$. Then, $\theta^{\prime}$ and $\phi^{\prime}$ are updated, as well as the weights and the sampled $z$. Then, the process is iterated until convergence.
2.2.1 Example: $\mathcal{L}(\phi,\theta)$ under normality
For illustration purposes, we develop below the case where $p_{\phi}$ and $p_{\theta}$ are normal distributions, similar to kingma2014vae [7]. As a reminder, any other distributions could be adopted, including non-continuous or non-numerical outcomes.
Encoder $p_{\phi}(z|x)$.
Let $m_{k}$ be the length of the latent variable $z_{k}$, $k=1,2$. We assume a normal model for $Z_{k}$ given $X_{k}=x_{k}$,
$Z_{k}|X_{k}=x_{k}\sim{\cal N}_{m_{k}}(g_{\phi_{k}}(x_{k}),\sigma_{k}^{2}J_{m_{k}}),$ (6)
where $J_{m_{k}}$ is the $m_{k}\times m_{k}$ identity matrix. In particular,
$\log p_{\phi}(z_{k}|x_{k})$ $=$ $-\frac{m_{k}}{2}\log 2\pi-\frac{m_{k}}{2}\log\sigma_{k}^{2}-\frac{1}{2\sigma_{k}^{2}}\sum_{j=1}^{m_{k}}\{z_{k,j}-g_{\phi_{k},j}(x_{k})\}^{2},\quad k=1,2.$ (7)
The mean $g_{\phi_{k}}(x_{k})$ can be any model, such as a neural network, with output of length $m_{k}$, $k=1,2$. The scale $\sigma_{k}$ can be fixed, computed via the weighted residuals, or learned through a separate set of models. It controls the amount of noise introduced in the latent dimension and is pivotal in determining the prediction interval width for $p(y|z)$. The code implementation allows for all three methods. This paper only presents the results for a fixed $\sigma_{k}$, as well as training a separate set of $g_{\phi_{k}}(x_{k})$ models to estimate varying $\sigma_{k}$s.
Decoder $p_{\theta}(y|z)$.
We assume a normal model for $Y$ given $Z=z$,
$Y|Z=z\sim\mathcal{N}(f_{\theta}(z),\sigma^{2}).$ (8)
This results in a log-likelihood contribution,
$\log p_{\theta}(y|z)=-\frac{1}{2}\log 2\pi-\frac{1}{2}\log\sigma^{2}-\frac{1}{2\sigma^{2}}\{y-f_{\theta}(z)\}^{2}.$ (9)
Again, the mean $f_{\theta}(z)$ can be any model, such as a neural network.
Summary.
Overall, the M-step is
$\phi_{k}^{*}=\arg\min_{\phi_{k}}\sum_{i,r}w_{i,r}\sum_{j=1}^{m_{k}}\{z_{k,i,r,j}-g_{\phi_{k},j}(x_{k,i,r})\}^{2},\quad k=1,2,$ (10)
$\theta^{*}=\arg\min_{\theta}\sum_{i,r}w_{i,r}\{y_{i}-f_{\theta}(z_{i,r})\}^{2},$ (11)
$(\sigma^{*})^{2}=\frac{1}{N}\sum_{i,r}w_{i,r}\{y_{i}-f_{\theta^{*}}(z_{i,r})\}^{2},$ (12)
2.3 Training
For efficiency purpose, the training set, $\{1,\ldots,N\}$, is segmented into batches $\{b_{1},\ldots,b_{L}\}$ on which the index $i$ runs (and thus the denominator of Eq. 12 must be adapted accordingly). The process iterates for each batch until the maximum number of steps is reached or an early stopping criterion is satisfied. The full details are given in Algorithm 1. The framework requires tuning hyper-parameters such as the number of MC samples $R$, the number of latent nodes $m_{k}$, and the standard deviation $\sigma_{k}$ of $Z_{k}$. Monitoring the point prediction on a hold-out validation is important to combat overfitting and terminate the training early with a patience hyper-parameter. Moreover, due to the generative nature of SEMF, the variation resulting from the initial random seed is measured in Subsection 4.1. Additionally, the model-specific hyper-parameters ($p_{\phi}$ and $p_{\theta}$) are also discussed in the same subsection.
Algorithm 1 SEMF Training: two input sources
1:$y,x_{1},x_{2},R$
$\theta,\phi_{1},\phi_{2}$
3:Initialize $\theta,\phi_{1},\phi_{2}$
4:Initialize $D_{y},D_{z_{1}},D_{z_{2}},D_{x},D_{i}$ to $\emptyset$
5:Split $I=\{1,\ldots,N\}$ into $L$ batches $\{b_{1},\ldots,b_{L}\}$
6:for $\ell=1$, …, $L$ do
7:   for all $i$ in $b_{\ell}$ do
8:      for $r=1$, …, $R$ do
9:         Simulate $z_{1,i,r}\sim p_{\phi_{1}}(\cdot|x_{1,i})$
10:         Simulate $z_{2,i,r}\sim p_{\phi_{2}}(\cdot|x_{2,i})$
11:         Set $z_{i,r}=[z_{1,i,r},z_{2,i,r}]$
12:      end for
13:      for $r=1$, …, $R$ do
14:         Compute
$w_{i,r}=\frac{p_{\theta}(y_{i}|z_{i,r})}{\sum_{r=1}^{R}p_{\theta}(y_{i}|z_{i,r})}$
15:         Update $D_{y}\leftarrow D_{y}\cup[y_{i}|z_{i,r}|w_{i,r}]$
16:         Update $D_{z_{2}}\leftarrow D_{z_{2}}\cup[z_{2,i,r}|x_{2,i}|w_{i,r}]$
17:         Update $D_{z_{1}}\leftarrow D_{z_{1}}\cup[z_{1,i,r}|x_{1,i}|w_{i,r}]$
18:      end for
19:      Update $\theta\leftarrow Q_{y}(\theta,D_{y})$
20:      Update $\phi_{1}\leftarrow Q_{1}(\phi_{1},D_{z_{1}})$
21:      Update $\phi_{2}\leftarrow Q_{2}(\phi_{2},D_{z_{2}})$
22:   end for
23:end for
24:Check convergence; Go to step 4 if not
2.4 Inference
The encoder-decoder structure of SEMF entails the simulations of $z_{r}$ during inference, as depicted in Figure 1. In theory, any inference on $y$ given $x$ can be performed for $\hat{y}$, for instance the mean value $\hat{y}=\frac{1}{R}\sum_{r=1}^{R}f_{\theta}(z_{r})$, where $z_{r}\sim p_{\phi}(z|x)$ (see Algorithm 2 for the simulation scheme). For prediction intervals, a second simulation step is used,
$z_{r}\sim p(z|x),\quad\hat{y}_{r,s}\sim p_{\theta}(y|z_{r}),\quad r,s=1,\ldots,R.$ (13)
Prediction interval on $y$ given $x$ at a given level $\alpha$ follows as
$PI=\text{quantile}\left(\{\hat{y}_{r,s}\};\frac{\alpha}{2},1-\frac{\alpha}{2}\right).$ (14)
In order to calibrate these intervals, we can use conformal prediction [15], which we review in Subsection 3.2 and subsequently explain in Subsection 4.1.
Figure 1: Inference procedure with the SEMF’s learnable parameters $\phi_{k}$ and $\theta$. Here, we illustrate the number of input sources $k$ as $k=1,2$
Algorithm 2 SEMF Simulation
1:$\theta^{*},\phi_{1}^{*},\phi_{2}^{*},x_{1},x_{2},R$
2:$z_{i,r}$
3:for $r=1,\ldots,R$ do
4:   Simulate $z_{1,i,r}\sim p_{\phi_{1}^{*}}(\cdot|x_{1,i})$
5:   Simulate $z_{2,i,r}\sim p_{\phi_{2}^{*}}(\cdot|x_{2,i})$
6:   Set $z_{i,r}=[z_{1,i,r},z_{2,i,r}]$
7:end for
3 Related Work
3.1 Latent Representation Learning
Latent-representation learning typically relies on the encoder-decoder architecture introduced in Section 2. Classic examples are Auto-Encoders (AEs) and VAEs: the encoder $g_{\phi}$ maps a sample input $x$ to a latent variable $z$, and the decoder $f_{\theta}$ reconstructs the input $\hat{x}=f_{\theta}(z)$. Training jointly optimizes the parameter vectors $\phi$ and $\theta$ by minimizing the reconstruction loss. Unlike AEs, which focus solely on reconstruction, VAEs introduce a variational objective that aims to model the data distribution through the marginal likelihood $p_{\theta}(x)$ while fitting approximate posterior distributions over the latent variables [7, 16]. However, directly maximizing this likelihood is difficult [16]. Thus, alternative methods exist, such as maximizing the evidence lower bound (ELBO), which provides guarantees on the log-likelihood [17].
For supervised and semi-supervised tasks, latent representation learning can include task-specific predictions [18]. More specifically, models such as AEs follow the classical encoder-decoder objective while training a predictor $h_{\psi}(z)$ through an additional layer or model to estimate the output $y$. This dual objective helps in learning more task-relevant embeddings [19, 20]. Semi-supervised VAEs are similar, with the distinction that they couple the reconstruction loss of the unlabeled data with a variational approximation of latent variables. This is effective even with sparse labels [21, 22].
The EM algorithm has already been used for supervised learning tasks using specific models [8, 9, 10], where the goal has been point prediction with GMMs. Similarly, the EM algorithm adapts well to minimal supervision [11] and using labeled and unlabeled data in semi-supervised settings for both single and multiple modalities [12, 13]. Our work differs by modifying using MC sampling to generate prediction intervals with any ML model and, in theory, under any distribution.
3.2 Prediction Intervals
Crucial for estimating uncertainty, prediction intervals provide a range for regression outcomes. Common approaches include Bayesian methods [23, 24, 25], ensemble techniques [26, 27, 28], and quantile regression [29, 30]. Quantile regression specifically utilizes the pinball loss function to target desired quantiles, proving effective even for non-parametric models [31] and asymmetric distributions [30]. This loss is valuable both for individual models and for refining quantile estimates when used with ensembles [32]—a loss that can also be learned jointly as demonstrated by Simultaneous Quantile Regression (SQR) [33]. SQR represents a powerful neural network approach that jointly estimates multiple quantiles while maintaining their proper ordering, making it particularly effective for constructing prediction intervals and a strong baseline for uncertainty quantification tasks.
Complementary to these approaches, conformal prediction (CP) offers a general framework to construct and calibrate prediction intervals for any model [34]. Assuming data exchangeability, CP adjusts intervals to ensure they meet a pre-specified coverage probability, thereby enhancing reliability in applications demanding rigorous uncertainty quantification. Recent advances in CP include its integration with ensemble models for time series [35, 36] and Bayesian approaches such as Conformal Bayesian model averaging (CBMA) aggregates conformity scores from several Bayesian models to obtain optimal prediction sets with guarantees [37]. Furthermore, several CP frameworks incorporate data-dependent score functions have recently been proposed to improve the efficiency and validity of intervals derived via quantile estimators [38], boosting methods (XGBoost) to refine conformity scores [39] and locally adaptive techniques to enhance conditional validity [40].
4 Experimental Setup
4.1 Models
Our baseline consists of quantile regression models based on eXtreme Gradient Boosting (XGBoost) [41], Extremely Randomized Trees (ET) [42], and neural networks using SQR [33], all summarized and depicted in Table 1. To ensure consistency in our experimental setup, we align the families and hyper-parameters of $p_{\phi}$ and $p_{\theta}$ with our baseline models. For example, in the case of XGBoost in SEMF, we use $K$ XGBoosts, $g_{\phi_{k}}(x_{k})$, one for each input $x_{k}$, $k=1,\ldots,K$, and one XGBoost for $f_{\theta}(z)$ with the same hyper-parameters. We refer to the SEMF’s adoption of these models as MultiXGBs, MultiETs, and MultiMLPs. When establishing prediction intervals, we conformalize our prediction intervals according to [43] at an uncertainty tolerance of 5% for both the baseline and SEMF (Eq. 14).
On the notable hyper-parameters, we target larger $\sigma_{k}$ values that introduce more noise and produce better intervals than point predictions. For several datasets, we set $\sigma_{k}$ to train_residual_models, which trains separate models to predict the scale of the latent variables individually and is more suitable to heteroscedastic noise patterns. This approach often yields substantially improved interval predictions by capturing local uncertainty structures, though at the cost of longer computation times due to the additional model training. In our experience, using train_residual_models for $\sigma_{k}$ is recommended as a starting point for new applications, with fixed values being a computationally efficient alternative once the behavior of the dataset is better understood. The optimal models are then trained and tested with five different seeds, and the results are averaged, and the variability of the results is presented. The point prediction, $\hat{y}$, uses the mean inferred values. Appendix A contains more details on the hyper-parameters for each SEMF model and dataset introduced below.
SEMF Base Model Interval Prediction Baseline
MultiXGBs XGBoost
Trees: 100, Maximum depth: 6, Early stopping steps: 10
Quantile XGBoost
Same as point prediction baseline, XGBoost
MultiETs Extremely Randomized Trees
Trees: 100, Maximum depth: 10
Quantile Extremely Randomized Trees
Same as point prediction baseline, Extremely Randomized Trees
MultiMLPs Deep Neural Network
Hidden layers: 2, Nodes per layer: 100, Activation functions: ReLU, Epochs: 1000 or 5000, Learning rate: 0.001, Batch training, Early stopping steps: 100
Simultaneous Quantile Regression
Same as point prediction baseline, Deep Neural Network
Table 1: SEMF models, baselines, and hyper-parameters.
4.2 Datasets
4.2.1 Simulations
In our experiments, we generate synthetic datasets with $10^{3}$ observations and $k=2$ predictors from a standard normal distribution according to
$y=f(x)+\epsilon,$ (15)
where $f(x)$ in a simple setup is defined as
$f(x)=\sum_{i=1}^{k}\cos(x_{i})$ (16)
to isolate noise effects, and in a more complex setup, we instead generate
$f(x)=\sum_{i=1}^{k}\Bigl{(}x_{i}^{2}+0.5\,\sin(3\,x_{i})\Bigr{)}$ (17)
to introduce nonlinearity and heteroscedasticity. $\epsilon$ is drawn from one of four distributions: $\mathcal{N}(0,0.5)$, $\mathcal{U}(-0.5,0.5)$, a centered log-normal, or $\text{Gumbel}(0,0.5)$. Figure 2 illustrates an example of Eq. 17 with $k=1$ (for simplicity) and 500 observations, where blue dots indicate predictions with 95% prediction intervals and the red curve denotes $f(x)$.
For our simulation experiments, the underlying data is generated from a fixed seed with variability only in the model seeds (Subsection 4.1). We evaluate all three model variants (MultiXGBs, MultiETs, and MultiMLPs) against standard quantile regression baselines. For MultiXGBs and MultiETs, we use $R=10$ with 10 nodes per latent dimension, while MultiMLPs use $R=5$ with 20 nodes per latent dimension. All models employ early stopping with appropriate patience settings to prevent overfitting on the synthetic data. The data scaling and splits are the same as the benchmark data below.
Figure 2: Prediction intervals under different noise distributions for a one-dimensional $x$ generated from 500 observations. Each panel shows predictions (blue dots) with 95% prediction intervals (shaded regions) for a model trained on data with (a) Normal, (b) Uniform, (c) Log-normal, and (d) Gumbel noise. The red curve denotes $f(x)$, the underlying (deterministic) function.
4.2.2 Benchmark data
We systematically curate a subset of datasets from the OpenML-CTR23 [44] benchmark suite to evaluate and carry out our experiments. Initially comprising 35 datasets, we apply an exclusion criteria to refine this collection to 11 datasets. The details and overview are in Appendix B. We remove duplicated rows from all the datasets and carry out the scaling of all predictors, including the outcome. The features of these datasets are then treated as separate inputs to SEMF. In all our datasets, 70% of the data is used to train all models, 15% as a hold-out validation set to monitor SEMF’s performance, and 15% to evaluate the models. To combat overfitting, baseline models that benefit from early stopping are allocated another 15% from the training data. Lastly, it is essential to note that all data in SEMF are processed batch-wise, without employing mini-batch training, to ensure consistency and stability in the training process.
4.3 Metrics
The evaluation of prediction intervals employs Prediction Interval Coverage Probability (PICP), Normalized Mean Prediction Interval Width (NMPIW), Continuous Ranked Probability Score (CRPS) and the quantile (pinball) loss. Detailed definitions and closed-form expressions for these metrics (with the CRPS computed under a uniformity assumption on the predictive distribution) are provided in Appendix C. In addition, we introduce the Coverage-Width Ratio (CWR)
$\text{CWR}=\frac{\text{PICP}}{\text{NMPIW}},$ (18)
which quantifies the trade-off between reliability and precision.
In our case, measuring the performance of SEMF over the baseline models is far more critical than reviewing absolute metrics in isolation. For any metric above, this is computed as
$\text{Metric}_{\Delta}(\%)=\left(\frac{\text{Metric}_{\text{SEMF}}-\text{Metric}_{\text{Baseline}}}{\text{Metric}_{\text{Baseline}}}\right)\times 100,$ (19)
on which we base our decisions for selecting the best hyper-parameters as explained in Subsection C.2.
5 Experiments
5.1 Simulated Data
Table 2 and Table 3 summarize the performance of our SEMF variants on synthetic datasets generated using two predictors. The PICP shows that on average, all models across all experiments were able to achieve at least the 95% desired coverage probability.
$\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP PICP
MultiXGBs
normal 18% [10:35] 16% [10:28] 6% [3:10] 11% [5:17] -1% [-3:0] 0.95$\pm$0.01
uniform 25% [-2:58] 19% [-5:39] 10% [-1:23] 18% [0:32] -1% [-5:2] 0.96$\pm$0.03
lognormal 28% [8:59] 21% [8:39] 13% [8:18] 7% [2:14] -1% [-3:0] 0.96$\pm$0.01
gumbel 10% [-5:20] 8% [-9:18] 8% [1:14] 8% [0:16] 0% [-3:4] 0.96$\pm$0.01
MultiETs
normal 15% [1:32] 14% [0:27] 5% [0:8] 8% [0:16] -2% [-4:1] 0.95$\pm$0.01
uniform 16% [3:35] 13% [-1:28] 6% [3:10] 13% [8:21] 0% [-3:4] 0.97$\pm$0.01
lognormal 5% [-14:18] 3% [-18:17] 9% [5:13] -3% [-10:2] 0% [-2:2] 0.97$\pm$0.01
gumbel 3% [-4:14] 3% [-7:14] 8% [6:11] 1% [-2:5] 0% [-2:2] 0.96$\pm$0.01
MultiMLPs
normal 3% [-2:9] 3% [-2:8] 1% [-2:4] 2% [-3:5] 0% [-1:0] 0.96$\pm$0.01
uniform 4% [-2:13] 4% [-2:10] 2% [-1:8] 7% [-5:23] 0% [-4:3] 0.97$\pm$0.02
lognormal 7% [-4:18] 6% [-5:16] 3% [-2:7] 4% [0:9] 0% [-1:1] 0.97$\pm$0.01
gumbel 5% [-3:13] 5% [-5:14] 0% [-2:3] 0% [-4:4] -1% [-2:2] 0.96$\pm$0.02
Table 2: Test results (95% prediction intervals) for 1000 observations generated using cosine $f(x)$ with 2 predictors, and an additive noise ($\epsilon$) belonging to one of the four distributions. Relative metrics are shown over 5 seeds as ‘mean [min:max]’, and absolute metrics as ‘mean $\pm$ std’. Average performance over the baseline is highlighted in bold.
$\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP PICP
MultiXGBs
normal 39% [28:65] 28% [22:41] 29% [26:34] 23% [16:35] 0% [-3:0] 0.95$\pm$0.01
uniform 65% [54:80] 39% [34:44] 41% [39:47] 25% [16:32] 0% [-1:1] 0.96$\pm$0.02
lognormal 0% [-13:7] -2% [-16:6] 20% [14:27] 4% [-4:16] 2% [0:5] 0.96$\pm$0.02
gumbel 23% [0:49] 17% [0:35] 20% [9:30] 16% [3:21] 0% [-3:4] 0.96$\pm$0.01
MultiETs
normal 23% [7:35] 19% [7:26] 14% [0:21] 12% [-4:23] -1% [-2:1] 0.95$\pm$0.02
uniform 40% [25:71] 28% [20:42] 24% [16:33] 12% [-4:21] 0% [-3:2] 0.97$\pm$0.02
lognormal 8% [-9:33] 5% [-13:27] 11% [0:16] 6% [-6:12] 1% [-2:3] 0.97$\pm$0.01
gumbel 29% [-3:68] 19% [-4:42] 18% [5:30] 15% [0:28] 0% [-2:1] 0.97$\pm$0.01
MultiMLPs
normal 5% [-14:24] 1% [-20:21] 7% [2:13] 13% [6:20] 2% [-2:4] 0.95$\pm$0.01
uniform 37% [15:66] 25% [13:41] 19% [13:27] 29% [9:40] 1% [-2:6] 0.97$\pm$0.02
lognormal 1% [-12:10] 1% [-15:11] 4% [2:8] -1% [-10:4] 0% [-2:1] 0.96$\pm$0.01
gumbel 15% [-6:38] 11% [-8:30] 10% [-3:23] 11% [-5:23] 0% [-3:2] 0.96$\pm$0.00
Table 3: Test results (95% prediction intervals) for 1000 observations generated using quadratic $f(x)$ with 2 predictors, and an additive noise ($\epsilon$) belonging to one of the four distributions. Relative metrics are shown over 5 seeds as ‘mean [min:max]’, and absolute metrics as ‘mean $\pm$ std’. Average performance over the baseline is highlighted in bold.
For the cosine experiments (Table 2), all three methods (MultiXGBs, MultiETs, and MultiMLPs) exhibit positive relative improvements over the baseline in terms of the CWR, NMPIW, CRPS, and pinball loss. For instance, under normal noise, the MultiXGBs variant improves the $\Delta$CWR by an average of 18% (ranging between 10% and 35%), while still preserving a PICP of approximately 0.95. Similar positive trends are seen across the uniform and lognormal noise conditions. Although under Gumbel noise the improvements are less pronounced, SEMF still delivers performance that is at least on par with the baseline.
For the quadratic experiments (Table 3), which represents a more complex heteroscedastic case, the improvements are even more substantial. Under normal noise, MultiXGBs achieve an average improvement of 39% in $\Delta$CWR, and, when the noise is uniform, gains can be as high as 65%. There is some variability—for example, under lognormal noise, the improvement in $\Delta$CWR for MultiXGBs is close to zero—but overall, the results indicate that SEMF produces tighter and more reliable prediction intervals compared to the baseline. Similar to the cosine case, the Gumbel noise setting shows the smallest relative gains; nevertheless, the performance under Gumbel noise remains competitive with the baseline.
5.2 Benchmark Data
We trained and tested 165 models corresponding to the three model types—MultiXGBs, MultiETs, and MultiMLPs—across 11 datasets, using five seeds for each combination. Table 4 presents the means and standard deviations for our metrics aggregated over the five seeds. Appendix E includes the results from each individual run.
Relative Metrics Absolute Metrics
Dataset $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
MultiXGBs
space_ga 11% [4:21] 9% [2:19] 11% [10:12] 9% [6:13] 0% [-2:2] 4.51$\pm$1.64 0.24$\pm$0.08 0.95$\pm$0.01
cpu_activity 5% [-3:11] 5% [-4:11] 16% [10:26] 15% [0:36] 0% [-1:1] 9.77$\pm$0.55 0.10$\pm$0.01 0.95$\pm$0.01
naval_propulsion_plant 163% [114:200] 61% [53:68] 72% [66:74] 45% [31:50] 0% [-3:2] 8.27$\pm$0.68 0.12$\pm$0.01 0.95$\pm$0.01
miami_housing 5% [-4:12] 5% [-4:10] 12% [8:17] -12% [-21:-5] 0% [-1:0] 8.75$\pm$0.36 0.11$\pm$0.00 0.95$\pm$0.00
kin8nm 18% [16:19] 16% [14:17] 23% [20:25] 8% [3:11] -1% [-1:-1] 2.31$\pm$0.04 0.41$\pm$0.01 0.94$\pm$0.00
concrete_compressive_strength 40% [24:73] 29% [20:45] 32% [27:35] 12% [4:25] -3% [-5:-1] 3.22$\pm$0.21 0.29$\pm$0.02 0.94$\pm$0.02
cars 32% [10:67] 24% [10:42] 27% [20:42] 17% [6:32] -1% [-3:0] 5.33$\pm$0.55 0.18$\pm$0.02 0.95$\pm$0.01
energy_efficiency 128% [72:230] 53% [39:68] 67% [59:81] 54% [43:73] 3% [-2:8] 13.95$\pm$2.46 0.07$\pm$0.01 0.96$\pm$0.00
california_housing 1% [-1:4] 1% [-1:4] 18% [15:23] -8% [-10:-3] 0% [-1:0] 2.24$\pm$0.08 0.42$\pm$0.02 0.95$\pm$0.01
airfoil_self_noise -8% [-32:6] -12% [-47:6] 2% [-36:20] -17% [-43:4] 0% [-1:0] 2.29$\pm$0.38 0.44$\pm$0.10 0.97$\pm$0.01
QSAR_fish_toxicity 19% [2:53] 14% [1:35] 13% [8:24] 15% [8:30] 0% [-1:2] 1.71$\pm$0.10 0.57$\pm$0.04 0.98$\pm$0.01
MultiETs
space_ga 5% [-2:15] 6% [-2:16] 8% [6:11] -6% [-9:-3] -1% [-3:0] 4.38$\pm$1.82 0.26$\pm$0.09 0.95$\pm$0.02
cpu_activity 2% [-1:5] 6% [2:10] 3% [-2:7] -18% [-22:-11] -4% [-5:-3] 8.25$\pm$0.23 0.11$\pm$0.00 0.95$\pm$0.01
naval_propulsion_plant 144% [118:160] 60% [56:63] 68% [64:72] 44% [40:52] -2% [-6:0] 3.65$\pm$0.25 0.26$\pm$0.02 0.95$\pm$0.01
miami_housing -9% [-15:-5] -10% [-19:-4] -3% [-7:0] -73% [-81:-67] 0% [0:1] 6.44$\pm$0.27 0.15$\pm$0.01 0.95$\pm$0.00
kin8nm 7% [5:10] 7% [4:8] 14% [12:17] -1% [-3:0] 0% [0:1] 2.13$\pm$0.05 0.45$\pm$0.01 0.95$\pm$0.01
concrete_compressive_strength 6% [-14:26] 4% [-20:23] 9% [4:12] -3% [-12:3] 0% [-3:3] 3.05$\pm$0.25 0.31$\pm$0.03 0.95$\pm$0.01
cars 15% [-25:54] 7% [-40:36] 6% [-8:20] 8% [-5:28] -1% [-3:4] 4.89$\pm$0.81 0.20$\pm$0.04 0.95$\pm$0.02
energy_efficiency 15% [1:26] 14% [1:23] 3% [-15:12] -11% [-80:25] -2% [-3:0] 15.93$\pm$2.57 0.06$\pm$0.01 0.95$\pm$0.03
california_housing 1% [-11:9] 2% [-11:11] 21% [16:25] -22% [-26:-17] -2% [-3:-1] 1.70$\pm$0.09 0.56$\pm$0.04 0.95$\pm$0.01
airfoil_self_noise -13% [-39:22] -20% [-66:18] -17% [-52:19] -26% [-61:18] -1% [-2:1] 2.25$\pm$0.46 0.45$\pm$0.10 0.97$\pm$0.01
QSAR_fish_toxicity -3% [-20:17] -6% [-26:18] -3% [-16:6] -8% [-19:2] -1% [-4:0] 1.71$\pm$0.22 0.58$\pm$0.09 0.97$\pm$0.01
MultiMLPs
space_ga 5% [-6:25] 6% [-4:21] 2% [-9:13] 0% [-13:8] -2% [-3:0] 5.10$\pm$1.78 0.21$\pm$0.07 0.94$\pm$0.02
cpu_activity -14% [-16:-9] -16% [-20:-11] -9% [-14:-1] -3% [-11:17] 0% [-1:1] 8.57$\pm$0.14 0.11$\pm$0.00 0.95$\pm$0.01
naval_propulsion_plant 23% [-2:84] 15% [-1:46] 17% [-3:46] -7% [-25:33] 0% [-1:1] 12.78$\pm$1.86 0.08$\pm$0.01 0.95$\pm$0.01
miami_housing -21% [-35:-13] -28% [-56:-15] -13% [-23:-9] -26% [-38:-11] 0% [-1:0] 9.06$\pm$1.06 0.11$\pm$0.02 0.95$\pm$0.01
kin8nm 9% [2:18] 8% [2:16] 8% [5:14] 8% [0:15] 0% [-2:1] 4.74$\pm$0.17 0.20$\pm$0.01 0.94$\pm$0.01
concrete_compressive_strength 31% [-5:73] 21% [-2:43] 16% [-2:35] 7% [-19:31] -2% [-4:2] 3.43$\pm$0.17 0.28$\pm$0.02 0.95$\pm$0.02
cars -12% [-44:21] -21% [-89:22] -12% [-23:5] -19% [-29:-13] 0% [-5:7] 5.40$\pm$1.25 0.19$\pm$0.04 0.96$\pm$0.03
energy_efficiency 87% [46:149] 45% [30:63] 42% [33:56] 23% [0:50] -1% [-6:5] 21.91$\pm$3.25 0.04$\pm$0.01 0.94$\pm$0.02
california_housing -11% [-14:-7] -12% [-16:-6] -2% [-5:0] -9% [-13:-3] 0% [-1:0] 2.13$\pm$0.08 0.45$\pm$0.02 0.95$\pm$0.01
airfoil_self_noise 42% [16:63] 28% [14:39] 30% [17:43] 23% [5:32] -1% [-2:1] 4.92$\pm$0.31 0.20$\pm$0.01 0.97$\pm$0.01
QSAR_fish_toxicity 8% [-15:36] 6% [-19:28] 1% [-9:13] 3% [-11:15] -1% [-3:1] 1.71$\pm$0.22 0.58$\pm$0.08 0.97$\pm$0.01
Table 4: Test results for all models with at 95% quantiles aggregated over five seeds. For relative metrics, values are shown as ‘mean[min:max]’, and absolute metrics as ‘mean $\pm$ std’. Performance over the baseline is highlighted in bold.
The results on OpenML-CTR23 datasets reveal distinct performance patterns across our model variants. MultiXGBs shows improvement over the baseline in most datasets, with notable results on naval_propulsion_plant ($\Delta$CWR: 163%, $\Delta$NMPIW: 61%) and energy_efficiency ($\Delta$CWR: 128%, $\Delta$NMPIW: 53%). For concrete_compressive_strength and cars, MultiXGBs yields moderate improvements in interval quality ($\Delta$CWR: 40% and 32%, respectively). On airfoil_self_noise, however, MultiXGBs performs slightly below the baseline. While maintaining PICP values comparable to the baseline (within ±3%), MultiXGBs generally produces narrower prediction intervals, resulting in higher CWR values.
MultiETs shows more variable performance. It achieves relative improvements on naval_propulsion_plant ($\Delta$CWR: 144%, $\Delta$NMPIW: 60%) and modest gains on energy_efficiency and cars, but underperforms on datasets like miami_housing, airfoil_self_noise, as well as QSAR_fish_toxicity. Interestingly, MultiETs often shows positive $\Delta$CRPS values even when the $\Delta$Pinball is negative, suggesting a trade-off between interval sharpness and quantile calibration.
MultiMLPs exhibits mixed results across datasets. It performs well on energy_efficiency ($\Delta$CWR: 87%, $\Delta$NMPIW: 45%) and airfoil_self_noise ($\Delta$CWR: 42%, $\Delta$NMPIW: 28%), but shows lower performance on cpu_activity, miami_housing, and california_housing. The absolute CWR for MultiMLPs on energy_efficiency (21.91) indicates efficient intervals that maintain the target coverage with minimal width.
Across all models, PICP values remain within the target range of approximately 0.95±0.03, indicating that SEMF maintains proper coverage while often reducing interval widths. These results suggest that SEMF’s benefits are most pronounced when applied to XGBoost, followed by the randomized trees and neural networks, the order depending on the metric.
5.3 Discussion
Our results indicate that overall, SEMF enhances uncertainty quantification in regression tasks, with varying effectiveness depending on the base model type and dataset characteristics. The framework often produces narrower prediction intervals while maintaining coverage targets, addressing the trade-off between interval width and reliability. The mechanism behind SEMF’s effectiveness appears to be model-dependent. For tree-based methods like XGBoost, which have limited capacity to model uncertainty directly, SEMF’s latent representation learning introduces a form of distributional modeling that enables more efficient uncertainty quantification. The iterative sampling and weighting procedure in the EM algorithm allows XGBoost to better capture the uncertainty structure in the data. This explains why MultiXGBs show the most consistent improvements across datasets.
Neural networks, which inherently have greater representational capacity, benefit differently from SEMF. The framework’s structured approach to uncertainty quantification through latent variables complements the neural network’s flexibility, particularly in datasets with complex feature interactions like energy_efficiency. However, this advantage is not universal across all datasets, suggesting that the interaction between SEMF’s latent structure and the neural network’s own representational capabilities may sometimes create redundancies. Even in datasets where SEMF shows more modest improvements (such as space_ga and kin8nm), it still performs at least on par with SQR, our strong neural network baseline for simultaneous quantile estimation. This performance floor across diverse datasets highlights SEMF’s robustness in different modeling scenarios.
The more variability in the performance of MultiETs compared to XGBoost can be attributed to two factors. First, the randomized splitting mechanism in ETs introduces inherent variability that may interfere with SEMF’s iterative refinement process. Second, ETs sometimes had less stable estimates for the weights (Eq. 5), affecting the quality of the latent representations learned during training. This hypothesis is supported by the higher standard deviations observed in MultiETs’ performance metrics compared to the other models.
Dataset characteristics also influence SEMF’s effectiveness. The framework shows particular strength on datasets like naval_propulsion_plant and energy_efficiency, which feature complex interactions among input variables that benefit from latent representation learning. Particularly for naval_propulsion_plant, the overall $y$ output is uniformly distributed, which the trees baselines fail to capture, but SEMF is more robust to this. Conversely, on simpler datasets or those with highly linear relationships, the additional complexity introduced by SEMF may not provide significant advantages over traditional quantile regression approaches. The relationship between CRPS and pinball Loss also provides insight into SEMF’s behavior. For several dataset-model combinations (e.g., MultiXGBs on miami_housing and california_housing), we observe positive $\Delta$CRPS alongside negative $\Delta$Pinball values. This divergence stems from fundamental differences in what these metrics evaluate. CRPS assesses the overall quality of the prediction interval, favoring narrow intervals that are well-centered around the true value. In contrast, pinball Loss measures the calibration of individual quantile predictions, penalizing miscalibrations at the tails more heavily. SEMF typically produces sharper intervals that may sacrifice some calibration at the exact quantile levels, resulting in better CRPS but potentially worse pinball loss compared to baseline models with wider but more conservatively calibrated intervals.
The stability of SEMF across different random seeds, evidenced by the relatively small standard deviations in metrics like PICP and NMPIW, contrasts with the higher variability in baseline models. This suggests that SEMF’s sampling-based approach leads to more consistent uncertainty estimates, an important consideration for applications where predictive stability is required. The tight clustering of PICP values around the target level (0.95) across diverse datasets demonstrates SEMF’s ability to maintain calibration while improving efficiency. Further, our experimental design focused on the conformalized performance of SEMF, with CP applied consistently across all models. The performance improvements observed with this standardized approach suggest that SEMF could benefit further from specialized calibration techniques designed to leverage its sampling-based uncertainty estimation. Future work might explore adaptive conformalization methods that account for the specific characteristics of SEMF’s predictive distributions.
6 Conclusion
This paper introduces the Supervised Expectation-Maximization Framework (SEMF), a novel model-agnostic approach for generating prediction intervals in datasets with any machine learning model. SEMF draws from the EM algorithm for supervised learning to devise latent representations that produce better prediction intervals than quantile regression. Our comprehensive evaluation demonstrates SEMF’s effectiveness in two complementary settings. First, controlled simulations across different noise distributions (normal, uniform, log-normal, and Gumbel) and data generating functions (cosine and quadratic with periodic components) show SEMF’s robustness to varying data characteristics, with particularly strong improvements for heteroscedastic data patterns. Second, a set of 165 experimental runs on 11 real-world benchmark datasets with three different model types confirmed that SEMF outperforms quantile regression in practical applications, particularly when using XGBoost, which intrinsically lacks latent representations. The framework’s ability to maintain consistent performance across different noise distributions while producing narrower intervals underscores its potential in various application domains and opens new avenues for further exploration of supervised latent representation learning and uncertainty estimation.
7 Limitations & Future Work
The primary limitation of this study was its reliance on the normality assumption, which may not fully capture the potential of SEMF across diverse data distributions. Furthermore, in our simulations, we show that our framework can learn non-normal patterns; however, further investigation examination of how SEMF under broader distributional parameters can be interesting. The computational complexity of the approach presents another significant challenge, as the current implementation can be optimized for large-scale applications. Additionally, while the CWR metric is valid, it implicitly assumes that a 1% drop in PICP equates to a 1% reduction in NMPIW, thus assuming a uniform distribution. Evaluating CWR under various distributional assumptions would provide a more comprehensive assessment of its implications.
Future work presents several intriguing avenues for exploration. A promising direction is the application of SEMF in multi-modal data settings, where the distinct $p_{\phi}$ components of the framework could be adapted to process diverse data types—from images and text to tabular datasets—enabling a more nuanced and powerful approach to integrating heterogeneous data sources. This capability positions the framework as a versatile tool for addressing missing data challenges across various domains and can also help expand it to discrete and multiple outputs. Another valuable area for development is the exploration of methods to capture and leverage dependencies among input features, which could improve the model’s predictive performance and provide deeper insights into the underlying data structure. These advancements can enhance the broader appeal of end-to-end approaches like SEMF in the ML community.
Appendix A Optimal set of hyper-parameters
Remark.
The datasets mentioned here have been explained in Table 6 and Appendix B.
The hyper-parameter tuning for SEMF is implemented and monitored using Weights & Biases [45]. A random search is done in the hyper-parameter space for a maximum of 500 iterations on all 11 datasets, focusing on tuning the models only on the the non-conformalized performance. Key hyper-parameters are varied across a predefined set to balance accuracy and computational efficiency. The following grid is used for hyper-parameter tuning: the number of importance sampling operations $R\in\{5,10,25,50,100\}$ (100 is omitted for MultiMLPS), nodes per latent dimension $m_{k}\in\{1,5,10,20,30\}$, and standard deviations $\sigma_{k}\in\{0.001,0.01,0.1,1.0\}$. For some datasets, the $\sigma_{k}$ parameter is set to train_residual_models. This indicates that a separate residual model is trained to predict the latent-variable scale, which can be particularly suitable in heteroscedastic scenarios where noise levels vary across instances. Early stopping steps (patience) are set to five or ten, and $R_{\text{infer}}$, the $R$ for inference as $R_{\text{infer}}$, is explored at [30, 50, 70]. For california_housing and cpu_activity, the $R_{\text{infer}}$ value is set to 30, while for all the other datasets, it is set to 50. We do this to ensure efficient computation, speed, and memory usage (especially for the GPU). Finally, the option to run the models in parallel must be consistently enabled. Table 5 shows the optimal set of the selected hyper-parameters.
Dataset R $m_{k}$ $\sigma_{k}$ Patience $R_{\mathrm{infer}}$
MultiXGBs
space_ga 10 30 train_residual_models 5 50
cpu_activity 5 30 1 5 30
naval_propulsion_plant 5 30 0.01 5 50
miami_housing 5 10 train_residual_models 5 50
kin8nm 5 30 train_residual_models 10 50
concrete_compressive_strength 25 30 train_residual_models 5 50
cars 50 10 train_residual_models 10 50
energy_efficiency 5 1 0.01 10 50
california_housing 5 10 0.1 5 30
airfoil_self_noise 25 1 train_residual_models 10 50
QSAR_fish_toxicity 50 30 1 5 50
MultiETs
space_ga 10 30 train_residual_models 10 50
cpu_activity 5 30 1 5 30
naval_propulsion_plant 5 30 0.01 5 50
miami_housing 10 10 train_residual_models 5 50
kin8nm 5 30 train_residual_models 10 50
concrete_compressive_strength 25 30 train_residual_models 10 50
cars 100 5 train_residual_models 10 50
energy_efficiency 5 1 0.01 10 50
california_housing 5 10 0.1 5 30
airfoil_self_noise 25 1 train_residual_models 10 50
QSAR_fish_toxicity 50 30 train_residual_models 10 50
MultiMLPs
space_ga 25 10 train_residual_models 10 50
cpu_activity 5 20 0.001 5 30
naval_propulsion_plant 5 20 0.001 5 50
miami_housing 5 20 train_residual_models 5 50
kin8nm 5 20 train_residual_models 5 50
concrete_compressive_strength 5 30 train_residual_models 10 50
cars 5 30 train_residual_models 5 50
energy_efficiency 50 30 0.1 10 50
california_housing 5 20 0.01 5 30
airfoil_self_noise 25 10 train_residual_models 10 50
QSAR_fish_toxicity 50 30 train_residual_models 5 50
Table 5: Hyper-parameters for MultiXGBs, MultiETs, and MultiMLPs.
MultiXGBs and MultiMLPs benefit from early stopping to reduce computation time. Similarly, the baseline models for these instances use the same hyper-parameters for early stopping. Further, the number of epochs in the case of MultiMLPs is set as 1000, except for energy_efficiency and QSAR_fish_toxicity, where this is changed to 5000. Any model-specific hyperparameter we did not specify in this paper remains at the implementation’s default value (e.g., the number of leaves in XGBoost from [41]). Along with the supplementary code, we provide three additional CSV files: one for the results and hyperparameters of all 165 runs and the other two for the optimal hyperparameters of SEMF models, both raw (directly from SEMF) and conformalized.
For training MultiXGBs and MultiETs, the computations are performed in parallel using CPU cores (Intel® Core i9-13900KF). Due to this, MultiETs are not fully deterministic and may exhibit slight variations between runs. For MultiMLPs, they are done on a GPU (NVIDIA® GeForce RTX 4090) which should give deterministic results across runs. All the computations are done on a machine with 32 GB of memory. The code provides further details on hardware and reproducibility.
Appendix B Datasets for tabular benchmark
OpenML-CTR23 [44] datasets are selected in the following manner. The first criterion is to exclude datasets exceeding 30,000 instances or 30 features to maintain computational tractability. Moreover, we exclude the moneyball data [46] to control for missing values and any datasets with non-numeric features, such as those with temporal or ordinal data not encoded numerically. We then categorize the datasets based on size: small for those with less than ten features, medium for 10 to 19 features, and large for 20 to 29 features. We apply a similar size classification based on the number of instances, considering datasets with more than 10,000 instances as large. To avoid computational constraints, we exclude datasets that were large in both features and instances, ensuring a varied yet manageable set for our experiments. This leads us to the final list of 11 datasets listed in Table 6.
Dataset Name N Samples N Features OpenML Data ID Y [Min:Max] Source
space_ga 3,107 7 45402 [-3.06:0.1] [47]
cpu_activity 8,192 22 44978 [0:99] [48]
naval_propulsion_plant 11,934 15 44969 [0.95:1.0] [49]
miami_housing 13,932 16 44983 [72,000:2,650,000] [50]
kin8nm 8,192 9 44980 [0.04:1.46] [51]
concrete_compressive_strength 1,030 9 44959 [2.33:82.6] [52]
cars 804 18 44994 [8,639:70,756] [53]
energy_efficiency 768 9 44960 [6.01:43.1] [54]
california_housing 20,640 9 44977 [14,999:500,001] [55]
airfoil_self_noise 1,503 6 44957 [103.38:140.98] [56]
QSAR_fish_toxicity 908 7 44970 [0.053:9.612] [57]
Table 6: Summary of benchmark tabular datasets retained from [44]
Appendix C Metrics for prediction intervals
C.1 Metrics’ definitions
The most common metrics for evaluating prediction intervals [58, 59, 60, 61] are:
  • Prediction Interval Coverage Probability (PICP): This metric assesses the proportion of times the true value of the target variable falls within the constructed prediction intervals. For a set of test examples ${(x_{1},y_{1}),\ldots,(x_{N},y_{N})}$, a given level of confidence $\alpha$, and their corresponding prediction intervals ${I_{1},\ldots,I_{N}}$, the PICP is calculated as:
    $\text{PICP}=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}(y_{i}\in[L_{i},U_{i}]),$ (20)
    where $U_{i}$ and $L_{i}$ are the upper and lower bounds of the predicted values for the $i$-th instance. $y_{i}$ is the actual value of the $i$-th test example, and $\mathds{1}$ is the indicator function, which equals 1 if $y_{i}$ is in the interval $[L_{i},U_{i}]$ and 0 otherwise. $0\leq\text{{PICP}}\leq 1$ where PICP closer to 1 and higher than the confidence level $\alpha$ is favored.
  • Mean Prediction Interval Width (MPIW): The average width is computed as
    $\text{MPIW}=\frac{1}{N}\sum_{i=1}^{N}(U_{i}-L_{i}),$ (21)
    which shows the sharpness or uncertainty, where $0\leq\text{{MPIW}}<\infty$ and MPIW close to 0 is preferred.
  • Normalized Mean Prediction Interval Width (NMPIW): Since MPIW varies by dataset, it can be normalized by the range of the target variable
    $\text{NMPIW}=\frac{\text{MPIW}}{\max(y)-\min(y)}$ (22)
    where $\max(y)$ and $\min(y)$ are the maximum and minimum values of the target variable, respectively. The interpretation remains the same as MPIW.
  • Continuous Ranked Probability Score (CRPS): We can approximate a full predictive cumulative distribution function by assuming that the forecast is uniformly distributed between its lower and upper bounds. Let $L$ and $U$ denote the lower and upper bounds (corresponding to the $\alpha/2$ and $21-\alpha/$ 2 quantiles, respectively). Then, for an observation $y$, we define
    $\text{CRPS}(y,L,U)=\begin{cases}L-y+\frac{U-L}{3},&y\leq L,\\ \frac{(y-L)^{3}+(U-y)^{3}}{3\,(U-L)^{2}},&y\in[L,U],\\ y-U+\frac{U-L}{3},&y\geq U.\end{cases}$ (23)
    This closed-form expression yields a lower CRPS when the forecast is both sharper and better calibrated.
  • Quantile (pinball) Loss: For a given quantile level $\tau\in(0,1)$ and prediction $q$, the pinball loss is defined as:
    $\ell_{\tau}(y,q)=(y-q)\left(\tau-\mathds{1}\{y\leq q\}\right),$ (24)
    where $\mathds{1}\{y\leq q\}$ equals 1 if $y\leq q$ and 0 otherwise. This loss penalizes underestimation and overestimation asymmetrically and is used to tune the quantile predictions.
C.2 Impact of relative metrics for modeling
As our primary focus is on interval prediction, configurations demonstrating the most significant improvements in $\Delta$CWR and $\Delta$PICP are prioritized when selecting the optimal hyper-parameters. Furthermore, both $\Delta$PICP and $\Delta$CWR must be positive, indicating that we must at least have the same reliability of the baseline (PICP) with better or same interval ratios (CWR). In instances where no configuration meets the initial improvement criteria for both metrics, we relax the requirement for positive $\Delta$PICP to accept values greater than -5% and subsequently -10%, allowing us to consider configurations where SEMF significantly improves CWR, even if the PICP improvement is less marked but remains within an acceptable range for drawing comparisons.
Appendix D Full synthetic experiments results
Relative Metrics Absolute Metrics
Seed $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
Normal
0 17% 15% 6% 11% 0% 1.78 0.54 0.97
10 11% 11% 3% 5% -1% 1.60 0.60 0.96
20 19% 18% 8% 14% -2% 1.89 0.50 0.95
30 10% 10% 5% 10% -1% 1.99 0.48 0.96
40 35% 28% 10% 17% -3% 2.49 0.37 0.93
Uniform
0 -2% -5% -1% 0% 2% 2.36 0.42 0.98
10 34% 26% 12% 26% -1% 2.12 0.47 0.99
20 9% 8% 5% 12% 1% 2.11 0.47 0.99
30 28% 26% 9% 19% -5% 3.59 0.25 0.91
40 58% 39% 23% 32% -3% 3.61 0.26 0.95
Lognormal
0 23% 19% 15% 9% 0% 2.08 0.46 0.96
10 59% 39% 18% 7% -3% 2.07 0.46 0.95
20 8% 9% 8% 2% -1% 1.96 0.48 0.95
30 9% 8% 9% 5% 0% 1.59 0.62 0.98
40 38% 28% 15% 14% 0% 2.30 0.42 0.97
Gumbel
0 4% 1% 7% 16% 4% 1.61 0.60 0.98
10 17% 18% 7% 4% -3% 1.78 0.53 0.94
20 20% 17% 14% 16% -1% 2.01 0.48 0.96
30 -5% -9% 1% 0% 4% 1.75 0.55 0.96
40 15% 14% 10% 5% -1% 2.02 0.47 0.95
Table 7: Test results for MultiXGBs, over 5 seeds, with cosine $f(x)$ using 2 predictors, organized by distribution and seed. Relative metrics are in bold when positive.
Relative Metrics Absolute Metrics
Seed $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
Normal
0 8% 8% 4% 0% -1% 1.98 0.48 0.96
10 17% 18% 6% 12% -3% 1.66 0.57 0.94
20 1% 0% 0% 2% 1% 1.81 0.54 0.97
30 15% 14% 8% 16% -1% 1.99 0.49 0.97
40 32% 27% 8% 9% -4% 2.46 0.38 0.94
Uniform
0 35% 28% 10% 21% -3% 2.80 0.35 0.97
10 3% -1% 3% 8% 4% 2.45 0.41 0.99
20 7% 5% 4% 11% 1% 2.58 0.37 0.96
30 9% 8% 6% 14% 0% 3.42 0.28 0.97
40 29% 25% 7% 12% -3% 3.63 0.26 0.96
Lognormal
0 -14% -18% 8% -10% 1% 1.85 0.52 0.96
10 18% 17% 10% -2% -2% 1.87 0.51 0.95
20 12% 12% 13% 0% -1% 2.03 0.47 0.96
30 -6% -8% 5% -4% 2% 1.74 0.56 0.98
40 14% 13% 11% 2% -2% 2.10 0.46 0.97
Gumbel
0 -4% -7% 9% 0% 2% 1.66 0.59 0.97
10 8% 7% 8% -2% 0% 1.72 0.56 0.97
20 -2% -2% 7% 5% 0% 1.90 0.50 0.96
30 2% 4% 6% -2% -2% 1.85 0.51 0.95
40 14% 14% 11% 5% -2% 1.96 0.50 0.97
Table 8: Test results for MultiETs, over 5 seeds, with cosine $f(x)$ using 2 predictors, organized by distribution and seed. Relative metrics are in bold when positive.
Relative Metrics Absolute Metrics
Seed $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
Normal
0 0% 1% -2% 0% -1% 1.98 0.48 0.95
10 9% 8% 4% 5% 0% 1.71 0.56 0.96
20 -2% -1% -1% -3% -1% 2.04 0.46 0.94
30 -1% -2% 1% 3% 0% 2.20 0.44 0.97
40 7% 7% 1% 5% 0% 2.48 0.39 0.96
Uniform
0 2% 2% 0% 8% 0% 2.98 0.33 0.98
10 6% 2% 5% 14% 3% 2.80 0.35 0.98
20 3% 7% 0% -5% -4% 3.02 0.31 0.94
30 -2% -2% -1% -4% 0% 3.85 0.25 0.96
40 13% 10% 8% 23% 1% 3.67 0.27 0.99
Lognormal
0 2% 3% 1% 4% 0% 1.71 0.57 0.98
10 18% 16% 7% 8% 0% 1.99 0.48 0.95
20 -4% -5% -2% 0% 1% 1.86 0.52 0.98
30 15% 13% 7% 9% -1% 1.74 0.56 0.98
40 4% 5% 1% 2% -1% 2.10 0.46 0.97
Gumbel
0 -3% -1% -2% -4% -1% 1.54 0.64 0.98
10 8% 8% 1% 4% -1% 1.88 0.51 0.95
20 8% 8% 3% 4% 0% 2.14 0.44 0.94
30 13% 14% 2% -4% -2% 1.87 0.51 0.95
40 -3% -5% -2% -2% 2% 1.76 0.56 0.98
Table 9: Test results for MultiMLPs, over 5 seeds, with cosine $f(x)$ using 2 predictors, organized by distribution and seed. Relative metrics are in bold when positive.
Relative Metrics Absolute Metrics
Seed $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
Normal
0 35% 26% 28% 16% 0% 5.80 0.17 0.96
10 65% 41% 34% 35% -3% 4.80 0.20 0.96
20 40% 29% 31% 16% 0% 4.41 0.21 0.94
30 30% 23% 26% 23% 0% 5.32 0.18 0.95
40 28% 22% 27% 28% 0% 5.83 0.16 0.95
Uniform
0 80% 44% 42% 30% 0% 10.12 0.09 0.94
10 54% 34% 47% 32% 1% 5.78 0.17 0.98
20 62% 38% 40% 31% 0% 5.62 0.17 0.96
30 62% 39% 39% 16% -1% 7.28 0.13 0.95
40 68% 41% 39% 18% -1% 7.93 0.12 0.97
Lognormal
0 2% 1% 20% 0% 1% 4.37 0.22 0.96
10 -2% -2% 14% -4% 1% 2.92 0.34 0.99
20 7% 2% 27% 6% 5% 3.81 0.25 0.94
30 -13% -16% 15% 3% 1% 3.28 0.29 0.96
40 7% 6% 23% 16% 0% 3.42 0.29 0.98
Gumbel
0 34% 27% 30% 20% -2% 5.03 0.19 0.96
10 49% 35% 24% 21% -3% 3.70 0.26 0.95
20 0% 0% 9% 3% 1% 2.81 0.35 0.99
30 11% 7% 17% 14% 4% 4.49 0.21 0.95
40 18% 15% 21% 20% 1% 4.37 0.22 0.97
Table 10: Test results for MultiXGBs, over 5 seeds, with quadratic $f(x)$ using 2 predictors, organized by distribution and seed. Relative metrics are in bold when positive.
Relative Metrics Absolute Metrics
Seed $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
Normal
0 7% 7% 0% -4% -1% 5.77 0.16 0.92
10 35% 26% 21% 23% 0% 4.66 0.21 0.98
20 29% 24% 15% 14% -2% 4.69 0.20 0.93
30 26% 21% 15% 13% -1% 4.91 0.19 0.95
40 21% 16% 18% 12% 1% 4.79 0.20 0.98
Uniform
0 28% 23% 16% -4% -1% 8.81 0.11 0.94
10 25% 20% 24% 21% 1% 4.95 0.20 0.99
20 47% 34% 25% 7% -3% 4.51 0.21 0.96
30 30% 22% 22% 15% 2% 5.85 0.17 0.98
40 71% 42% 33% 21% 0% 7.34 0.13 0.97
Lognormal
0 -9% -13% 0% -6% 3% 4.34 0.22 0.96
10 -1% -2% 13% 12% 1% 3.36 0.29 0.98
20 33% 27% 16% 10% -2% 4.09 0.23 0.95
30 8% 8% 11% 9% 0% 3.29 0.30 0.97
40 10% 8% 15% 9% 1% 3.78 0.26 0.97
Gumbel
0 5% 4% 10% 7% 1% 4.67 0.21 0.98
10 42% 30% 27% 28% -1% 4.55 0.21 0.94
20 33% 25% 16% 15% -1% 3.44 0.28 0.97
30 -3% -4% 5% 0% 1% 3.86 0.25 0.97
40 68% 42% 30% 24% -2% 3.97 0.24 0.97
Table 11: Test results for MultiETs, over 5 seeds, with quadratic $f(x)$ using 2 predictors, organized by distribution and seed. Relative metrics are in bold when positive.
Relative Metrics Absolute Metrics
Seed $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
Normal
0 -14% -20% 3% 6% 3% 6.63 0.15 0.97
10 24% 21% 13% 19% -2% 6.07 0.16 0.96
20 0% -4% 2% 6% 3% 5.81 0.16 0.93
30 -1% -5% 7% 17% 4% 6.13 0.16 0.95
40 15% 14% 9% 20% 0% 7.47 0.13 0.95
Uniform
0 42% 26% 25% 36% 6% 15.86 0.06 0.97
10 27% 20% 14% 30% 2% 9.42 0.11 1.00
20 66% 41% 27% 29% -2% 10.02 0.10 0.95
30 34% 26% 13% 9% 0% 12.50 0.08 0.96
40 15% 13% 16% 40% -1% 13.66 0.07 0.97
Lognormal
0 1% 2% 4% 0% -1% 5.41 0.18 0.97
10 3% 2% 8% 0% 1% 5.23 0.18 0.96
20 10% 11% 3% -10% -2% 4.79 0.20 0.95
30 -12% -15% 4% 0% 1% 3.86 0.25 0.97
40 4% 3% 2% 4% 0% 4.88 0.20 0.97
Gumbel
0 38% 30% 19% 18% -3% 6.90 0.14 0.96
10 32% 25% 23% 23% -1% 4.56 0.21 0.96
20 -6% -8% -3% -5% 2% 4.10 0.23 0.96
30 13% 12% 3% 8% -1% 5.40 0.18 0.96
40 -3% -5% 10% 10% 1% 5.14 0.19 0.96
Table 12: Test results for MultiMLPs, over 5 seeds, with quadratic $f(x)$ using 2 predictors, organized by distribution and seed. Relative metrics are in bold when positive.
Appendix E Full conformalized results on benchmark datasets
Relative Metrics Absolute Metrics
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball CWR NMPIW PICP
space_ga 9% 0% 9% 12% 10% 6.28 0.15 0.95
space_ga 5% 2% 2% 11% 13% 2.95 0.32 0.95
space_ga 4% 1% 2% 10% 9% 3.23 0.30 0.97
space_ga 21% -2% 19% 11% 9% 3.37 0.28 0.96
space_ga 16% -2% 15% 11% 6% 6.74 0.14 0.94
cpu_activity 7% 0% 7% 17% 10% 9.84 0.10 0.96
cpu_activity -3% 1% -4% 10% 0% 8.69 0.11 0.96
cpu_activity 11% -1% 11% 14% 10% 10.20 0.09 0.94
cpu_activity 11% 1% 10% 26% 36% 10.14 0.09 0.95
cpu_activity 1% 1% 0% 13% 18% 9.99 0.10 0.95
naval_propulsion_plant 200% -3% 68% 73% 43% 9.19 0.10 0.92
naval_propulsion_plant 114% 0% 53% 66% 31% 7.21 0.13 0.95
naval_propulsion_plant 159% 1% 61% 74% 50% 7.97 0.12 0.96
naval_propulsion_plant 187% 0% 65% 74% 50% 8.73 0.11 0.95
naval_propulsion_plant 156% 2% 60% 71% 50% 8.24 0.12 0.96
miami_housing 9% -1% 9% 14% -5% 9.08 0.10 0.95
miami_housing -4% 0% -4% 8% -12% 8.28 0.12 0.96
miami_housing 2% 0% 3% 9% -19% 8.49 0.11 0.94
miami_housing 12% 0% 10% 17% -5% 9.23 0.10 0.95
miami_housing 8% -1% 9% 11% -21% 8.70 0.11 0.95
kin8nm 16% -1% 15% 22% 3% 2.29 0.41 0.94
kin8nm 19% -1% 16% 25% 11% 2.31 0.41 0.94
kin8nm 16% -1% 14% 20% 6% 2.27 0.42 0.95
kin8nm 19% -1% 17% 22% 8% 2.29 0.41 0.95
kin8nm 19% -1% 17% 24% 11% 2.40 0.40 0.95
concrete_compressive_strength 31% -1% 24% 35% 25% 2.95 0.33 0.96
concrete_compressive_strength 73% -5% 45% 34% 14% 3.60 0.26 0.92
concrete_compressive_strength 27% -3% 24% 29% 6% 3.14 0.29 0.92
concrete_compressive_strength 47% -2% 34% 33% 9% 3.14 0.30 0.94
concrete_compressive_strength 24% -1% 20% 27% 4% 3.25 0.29 0.93
cars 32% -2% 25% 25% 6% 5.78 0.17 0.95
cars 10% -1% 10% 20% 18% 5.02 0.19 0.95
cars 27% 0% 22% 26% 17% 5.08 0.19 0.95
cars 67% -3% 42% 42% 32% 6.15 0.15 0.94
cars 22% -2% 20% 23% 15% 4.63 0.21 0.97
energy_efficiency 114% 1% 52% 66% 43% 15.29 0.06 0.96
energy_efficiency 112% 0% 53% 67% 50% 14.18 0.07 0.95
energy_efficiency 109% -2% 53% 64% 44% 10.78 0.09 0.96
energy_efficiency 230% 8% 68% 81% 73% 17.66 0.05 0.96
energy_efficiency 72% 5% 39% 59% 58% 11.84 0.08 0.95
california_housing 3% 0% 3% 23% -3% 2.09 0.46 0.96
california_housing 4% -1% 4% 18% -10% 2.27 0.42 0.94
california_housing 2% 0% 1% 21% -7% 2.29 0.41 0.95
california_housing 0% 0% 0% 15% -9% 2.27 0.41 0.94
california_housing -1% 0% -1% 15% -10% 2.28 0.42 0.95
airfoil_self_noise -5% 0% -5% 9% -19% 2.60 0.37 0.96
airfoil_self_noise 6% 0% 6% 20% 4% 2.56 0.38 0.98
airfoil_self_noise -32% 0% -47% -36% -43% 1.57 0.63 0.99
airfoil_self_noise -4% 0% -5% 8% -10% 2.26 0.43 0.97
airfoil_self_noise -8% -1% -7% 9% -15% 2.43 0.39 0.95
QSAR_fish_toxicity 2% 0% 1% 8% 10% 1.80 0.54 0.97
QSAR_fish_toxicity 11% -1% 10% 8% 12% 1.63 0.59 0.96
QSAR_fish_toxicity 4% 2% 3% 9% 8% 1.80 0.54 0.97
QSAR_fish_toxicity 26% 0% 20% 17% 18% 1.55 0.64 0.99
QSAR_fish_toxicity 53% -1% 35% 24% 30% 1.73 0.57 0.98
Table 13: Test results for MultiXGBs with complete data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
Relative Metrics Absolute Metrics
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball CWR NMPIW PICP
space_ga 4% -2% 6% 6% -9% 5.86 0.16 0.95
space_ga -2% 0% -2% 7% -3% 2.66 0.36 0.96
space_ga 10% -1% 10% 7% -5% 3.19 0.30 0.97
space_ga -2% 0% -2% 7% -5% 2.99 0.33 0.97
space_ga 15% -3% 16% 11% -9% 7.22 0.13 0.91
cpu_activity 5% -5% 10% 2% -22% 8.69 0.11 0.94
cpu_activity 1% -3% 4% 7% -11% 8.05 0.12 0.96
cpu_activity 3% -5% 8% 5% -11% 8.26 0.11 0.94
cpu_activity -1% -4% 2% -2% -22% 8.14 0.12 0.95
cpu_activity 3% -4% 7% 2% -22% 8.11 0.12 0.95
naval_propulsion_plant 154% -6% 63% 69% 40% 3.55 0.27 0.94
naval_propulsion_plant 153% 0% 60% 68% 46% 3.95 0.24 0.95
naval_propulsion_plant 160% 0% 61% 72% 52% 3.86 0.25 0.96
naval_propulsion_plant 136% 0% 58% 66% 41% 3.67 0.26 0.95
naval_propulsion_plant 118% -4% 56% 64% 41% 3.23 0.29 0.95
miami_housing -13% 1% -16% -4% -75% 6.25 0.15 0.96
miami_housing -6% 0% -6% 0% -67% 6.71 0.14 0.95
miami_housing -5% 1% -6% 0% -67% 6.47 0.15 0.95
miami_housing -5% 0% -4% -1% -75% 6.75 0.14 0.95
miami_housing -15% 1% -19% -7% -81% 6.03 0.16 0.95
kin8nm 8% 0% 7% 15% 0% 2.14 0.44 0.95
kin8nm 7% 0% 7% 13% -3% 2.11 0.45 0.95
kin8nm 7% 0% 7% 14% 0% 2.10 0.46 0.96
kin8nm 5% 1% 4% 12% 0% 2.08 0.46 0.95
kin8nm 10% 1% 8% 17% 0% 2.22 0.43 0.95
concrete_compressive_strength 26% -3% 23% 12% 0% 3.27 0.29 0.94
concrete_compressive_strength 5% 0% 5% 8% 0% 2.97 0.33 0.97
concrete_compressive_strength 13% 3% 9% 12% 3% 3.39 0.28 0.93
concrete_compressive_strength -14% 2% -20% 4% -12% 2.73 0.35 0.95
concrete_compressive_strength 3% 0% 3% 9% -4% 2.87 0.34 0.97
cars 12% -1% 12% 3% 0% 4.95 0.19 0.96
cars 2% -2% 4% -1% 0% 4.97 0.19 0.96
cars 29% -3% 24% 15% 15% 3.64 0.27 0.97
cars -25% 4% -40% -8% -5% 6.19 0.15 0.92
cars 54% -2% 36% 20% 28% 4.70 0.21 0.97
energy_efficiency 26% -3% 23% 12% 0% 15.20 0.06 0.96
energy_efficiency 16% -3% 17% 6% 0% 16.14 0.06 0.94
energy_efficiency 24% 0% 20% 9% 25% 14.43 0.07 0.99
energy_efficiency 8% 0% 8% 4% 0% 20.70 0.04 0.91
energy_efficiency 1% -1% 1% -15% -80% 13.19 0.07 0.96
california_housing -$-1$1% 16% -26% 1.55 0.62 0.96
california_housing -1% -2% 1% 20% -24% 1.68 0.57 0.95
california_housing 2% -2% 4% 24% -17% 1.70 0.56 0.95
california_housing 4% -3% 7% 20% -25% 1.76 0.53 0.93
california_housing 9% -3% 11% 25% -17% 1.83 0.52 0.94
airfoil_self_noise -21% -1% -24% -26% -40% 2.15 0.45 0.97
airfoil_self_noise 22% 0% 18% 19% 18% 2.91 0.34 0.98
airfoil_self_noise -20% -2% -22% -21% -36% 2.17 0.44 0.96
airfoil_self_noise -39% 1% -66% -52% -61% 1.52 0.64 0.98
airfoil_self_noise -6% -1% -5% -3% -12% 2.49 0.39 0.97
QSAR_fish_toxicity -13% 0% -14% -4% -10% 1.69 0.57 0.97
QSAR_fish_toxicity -20% 0% -26% -16% -19% 1.36 0.72 0.98
QSAR_fish_toxicity 17% -4% 18% 6% -2% 2.00 0.47 0.94
QSAR_fish_toxicity 17% -2% 16% 4% 2% 1.87 0.52 0.97
QSAR_fish_toxicity -18% 0% -22% -7% -14% 1.60 0.61 0.97
Table 14: Test results for MultiETs with complete data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
Relative Metrics Absolute Metrics
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball CWR NMPIW PICP
space_ga 5% -3% 8% 0% -8% 7.12 0.13 0.92
space_ga -6% -2% -4% -9% -13% 3.01 0.32 0.95
space_ga 3% 0% 3% 2% 3% 3.60 0.27 0.97
space_ga 25% -2% 21% 13% 8% 4.46 0.21 0.93
space_ga 0% 0% -1% 3% 8% 7.29 0.13 0.94
cpu_activity -16% -1% -18% -14% -10% 8.51 0.11 0.95
cpu_activity -9% 0% -11% -9% -11% 8.44 0.11 0.96
cpu_activity -14% 0% -16% -11% -11% 8.43 0.11 0.96
cpu_activity -16% 1% -20% -1% 17% 8.73 0.11 0.95
cpu_activity -14% 0% -16% -8% 0% 8.75 0.11 0.95
naval_propulsion_plant 84% -1% 46% 46% 33% 15.95 0.06 0.94
naval_propulsion_plant 26% 0% 20% 27% 0% 12.02 0.08 0.95
naval_propulsion_plant -2% -1% -1% -3% -25% 13.52 0.07 0.94
naval_propulsion_plant 8% 1% 7% 14% -25% 11.94 0.08 0.96
naval_propulsion_plant 1% 0% 1% 2% -20% 10.46 0.09 0.95
miami_housing -19% -1% -23% -13% -35% 9.00 0.11 0.95
miami_housing -35% 0% -56% -23% -38% 7.03 0.14 0.96
miami_housing -13% 0% -15% -9% -33% 9.98 0.09 0.94
miami_housing -20% 0% -25% -9% -11% 9.56 0.10 0.94
miami_housing -17% 0% -21% -9% -12% 9.71 0.10 0.95
kin8nm 17% 0% 14% 14% 15% 4.50 0.21 0.96
kin8nm 4% 1% 3% 5% 6% 4.86 0.20 0.95
kin8nm 18% -2% 16% 11% 11% 4.97 0.19 0.93
kin8nm 2% 0% 2% 5% 6% 4.59 0.21 0.95
kin8nm 3% 0% 3% 5% 0% 4.80 0.20 0.93
concrete_compressive_strength 54% -4% 37% 26% 21% 3.33 0.29 0.96
concrete_compressive_strength -5% -2% -2% -2% -19% 3.15 0.30 0.94
concrete_compressive_strength 13% -3% 14% 8% -12% 3.51 0.26 0.92
concrete_compressive_strength 73% -2% 43% 35% 31% 3.47 0.28 0.97
concrete_compressive_strength 19% 2% 14% 17% 15% 3.67 0.26 0.95
cars -11% -2% -11% -13% -29% 5.13 0.18 0.94
cars -7% -1% -7% -9% -13% 5.21 0.19 0.97
cars -18% 0% -22% -17% -19% 4.00 0.24 0.98
cars 21% -5% 22% 5% -21% 7.75 0.12 0.92
cars -44% 7% -89% -23% -14% 4.89 0.20 0.99
energy_efficiency 102% 5% 48% 56% 50% 19.46 0.05 0.96
energy_efficiency 74% -1% 42% 40% 40% 19.62 0.05 0.95
energy_efficiency 65% -3% 41% 36% 25% 18.91 0.05 0.96
energy_efficiency 46% 3% 30% 33% 0% 26.91 0.03 0.90
energy_efficiency 149% -6% 63% 42% 0% 24.65 0.04 0.92
california_housing -10% 0% -11% 0% -3% 1.99 0.48 0.96
california_housing -14% 0% -16% -5% -13% 2.12 0.45 0.94
california_housing -11% 0% -12% -1% -9% 2.12 0.45 0.95
california_housing -7% -1% -6% 0% -9% 2.22 0.42 0.94
california_housing -12% 0% -14% -2% -9% 2.20 0.43 0.95
airfoil_self_noise 60% -1% 38% 33% 30% 5.45 0.18 0.97
airfoil_self_noise 55% -1% 36% 37% 27% 4.83 0.20 0.97
airfoil_self_noise 63% 0% 39% 43% 32% 4.98 0.20 0.98
airfoil_self_noise 17% 1% 14% 20% 23% 4.84 0.20 0.95
airfoil_self_noise 16% -2% 15% 17% 5% 4.48 0.21 0.95
QSAR_fish_toxicity 21% -3% 19% 4% 6% 1.85 0.52 0.96
QSAR_fish_toxicity 2% 1% 1% 1% 2% 1.57 0.62 0.98
QSAR_fish_toxicity -2% 0% -2% -3% 2% 1.67 0.58 0.97
QSAR_fish_toxicity 36% -3% 28% 13% 15% 2.04 0.47 0.96
QSAR_fish_toxicity -15% 1% -19% -9% -11% 1.43 0.69 0.98
Table 15: Test results for MultiMLPs with complete data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.